OpenStudy (anonymous):

Lines k and p are perpendicular, neither is vertical and p passes through the origin. Which is greater? A. The product of the slopes of lines k and p B. The product of the y-intercepts of lines k and p

OpenStudy (irishboy123):

this is a chain of thought: \(y_k = m_k x + c_k\) \(y_p = m_p x + c_p\) [A]: "p passes through the origin. " \(c_p = 0\) [B]: "Lines k and p are perpendicular" \(m_k \times m_p = -1\) "Which is greater?" "A. The product of the slopes of lines k and p " \(m_k \times m_p = -1\) B. The product of the y-intercepts of lines k and p \(c_p = 0\) and "neither is vertical" \(\implies c_k \times c_p = 0\) does that help you or hinder you??!!

OpenStudy (anonymous):

OMFG

OpenStudy (anonymous):

HINDER HINDER DEFINITELY HINDER

OpenStudy (irishboy123):

soz to shorten: line p goes through origin so its intercept is zero. the lines, k & p, are perpendicular .... so the products of the slopes is -1.

OpenStudy (anonymous):

which intercept? y or x?

OpenStudy (irishboy123):

it goes through the Origin. it is therefore y = mx. it hits both axes at (0,0)

OpenStudy (anonymous):

hmmmmmmmmm

OpenStudy (irishboy123):

soz @yomamabf , haven't helped, have i ?!?! but: bed, tired. hope you get there :p